OpenStudy (anonymous):

Use substitution to solve the following systems of equations: 4y+x=11 3x+2y+13

OpenStudy (anonymous):

x = 11-4y put this in eq2

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

3(x = 11-4y) + 2y = 13 :-/

OpenStudy (anonymous):

3( 11-4y) + 2y = 13

OpenStudy (anonymous):

so thats the answer.?

OpenStudy (anonymous):

um idk...find out